Prove that f:R→ R defined by f(x)=x3 -5 is a bijection
1
Expert's answer
2020-09-24T16:49:51-0400

"f: \real \to \real" is a bijection if and only if f(x) is an injection and is Surjective.

Injection

If "f(x) = f(y) \implies x =y" then f is an injection.


"x^3-5=y^3-5""x^3=y^3""x=y"

Therefore, f is an injection.

Surjection:

"f: A \to B" is Surjective if for all "y \in B \exists x \in A: F(x) =y"


"y=x^3-5""x={(y+5)}^{1 \over 3}"

"f(x) ={({(y+5)^{1 \over 3}})^3}-5""=y+5-5""=y"

Therefore f is Surjective

Thus, since "f: \real \to \real" is both injective and Surjective, it is a bijection